40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Fri Nov 4 23:34:28 2016, awade, DailyProgress, Electronics Equipment, Trouble shooting in the north path 100_0258.JPG100_0261.JPG
    Reply  Tue Nov 8 14:01:11 2016, awade, DailyProgress, Electronics Equipment, Trouble shooting in the north path 2016-11-08_11.35.34.jpg
       Reply  Wed Nov 16 12:46:59 2016, awade, DailyProgress, Electronics Equipment, Trouble shooting in the north path and Murphy's law 
          Reply  Wed Dec 7 18:59:46 2016, awade, DailyProgress, Electronics Equipment, FSS monitor daughter board damage and Murphy's law D040424-B.pdf
             Reply  Wed Dec 7 21:14:07 2016, Koji, DailyProgress, Electronics Equipment, FSS monitor daughter board damage and Murphy's law 
                Reply  Thu Dec 8 17:37:32 2016, awade, DailyProgress, Electronics Equipment, FSS monitor daughter board fixed 
       Reply  Thu Jan 5 15:50:55 2017, awade, DailyProgress, Electronics Equipment, Trouble shooting in the north path 2017-01-05_15.17.55.jpg
Message ID: 1764     Entry time: Tue Nov 8 14:01:11 2016     In reply to: 1759     Reply to this: 1770   1803
Author: awade 
Type: DailyProgress 
Category: Electronics Equipment 
Subject: Trouble shooting in the north path 

Now with a useable error signal for the north path, I've been trying to track down the source of ringing in the north FFS control loop. Initially I though it was too much gain, but it turns out to be related to offset.  With the current settings the fast monitor output is showing ringing at ~1 MHz.  It turns out that this is because the offset is set to low, as it rails it set of this oscillation (pictured below) that persists until I reset the lock. Vpp of this oscillation is 1.08 V (min=-1.78 V, max=-0.680V).  The upper rail (at 1.52 V) didn't have this oscillation.

I blocked the lasers, turned off the Marconi output and also added terminators to the test input on the RF stage (I don't know why this wasn't done before).  Watching the voltage at the fast monitor output I adjusted the north offset until the fast voltage was zeroed.  It was extremely sensitive at the zero crossing point. Even when I left the knob untouched it was flickering up and down.  I tried the same thing on the south path with identical gain settings.  The offset smoothly tuned from one rail to the other (only going into oscillations at the lower rail) and was stable. It didn't have any of the flickering level of the north path. To clarify the north path pretty much flipped from one rail to the other with very little adjustment, the south path was much less sensitive.  I'm not sure what is going on here.  Possibly a issue with grounding? Or something about the stages of gain has been altered in the north TTFFS box? Something is broken?

 

Maybe when I relock the loop the OLG will suppress this residual noise. But the sensitivity to offset knob tuning my mean that it will drift back to generating this 1 MHz oscillation, which is above the UGF of the loop.  I think this is something that needs to be solved before a sensible beat note can be detected.

--

 

In another unrelated problem I noticed that the ±24 V power supplies were flickering on the edge of current limit mode.  We are pulling almost 1 amp @ 24V to run all the electronics in the rack.  Seem surprisingly high.  I turned up the current limit, the +24 V line is now giving full voltage as it should.

Quote:

I have been going through to optimize in the North path today. In the course of rearrangements of some electronics I turned off the power supply (Tenma 72-7245) to the north EOM 14.75 MHz amplifier to the broadband phase modulator.  The voltage settings were lost, so I though I might replace the Tenma with one of the HP power supplies.  Don't know if the Tenmas are particularly lorded for their low noise performance and the HPs are usually ok (although I haven't ever measured this).  

In the process I've broken off one of the banana plug pins on the EOM driver, a new one will need soldering on.  Also, the driver is apparently wrapped in a cardboard box. I'd say this should be replaced too.  We're definitely out of plastic jiffy boxes in the westbridge electronics lab, I might order some more of those. For my reference the dimensions of the board and its components are 50.75x30.5x18 mm.

North path locking is down until at least tomorrow while I do this small fix.

 

 

Attachment 1: 2016-11-08_11.35.34.jpg  2.372 MB  | Hide | Hide all
2016-11-08_11.35.34.jpg
ELOG V3.1.3-